Thinfilms morning report 2/24/05
jhaydon at snf.stanford.edu
Thu Feb 24 11:07:47 PST 2005
Thinfilms Morning Report 2/24/2005
Innotec; ion gauge controller not working I should have the
parts tomorrow innotec now has a leak that I will not be able to check
out until next week.
STS Dep: ok
Implanter: noisy signal still but ion gauge problem is fixed
if you look at the end station ion gauge I will try to get to the noise
problem next week
DI water: ok
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the thinfilms